OpenStudy (anonymous):

(-6) with exponent of 12 times (-6) with exponent of 5 times (-6) with exponent of 2?

OpenStudy (anonymous):

Hi, welcome to openstudy:) When you multiply a power with the same base, you just add the exponenets together. In this case: \[-6^{12} \times -6^5 \times -6^2\] you would add 12 by 5 by 2: \[(-6)^{12+5+2}\] Can you do the rest from here?

OpenStudy (anonymous):

Can you add the exponents?

OpenStudy (anonymous):

yes but the -6's are all in perentheses does that mean anything ?

OpenStudy (anonymous):

Not really. You would get \[(-6)^{19}\] You see, it's okay if it's a negative, because all three of them are negative. They have to be exactly the same for you to be able to add the exponents like that :) Does that help?

OpenStudy (anonymous):

okay and if they weren't exactly the same what would u do? like say one of the -6's is a 4

OpenStudy (anonymous):

The parentheses just indicate that the negative is part of the 6, so that it isn't confused with a subtraction sign. If one of the -6's was a 4, lets say \[(−6)^{12} × (−6)^{5} × (4)^2\] Then it would be \[(−6)^{12+5} × (4)^2\] or just \[(−6)^{17} × (4)^2\] To simplify further, you add the powers together and multiply the bases, so \[(−6)^{17} × (4)^2 = -24^{19} \]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yes? No? I hope this helped :)

OpenStudy (anonymous):

ohhhh okay